# Automatic Fact-Checking with Frame-Semantics

### v1

@@ Abstract
We propose a novel paradigm for automatic fact-checking that leverages frame semantics to enhance the structured understanding of claims, addressing the challenges posed by misinformation in today’s information ecosystem. To support this approach, we introduce a pilot dataset of real-world claims extracted from PolitiFact, specifically annotated for large-scale structured data. This dataset underpins two case studies: the first investigates voting-related claims using the Vote semantic frame, while the second explores various semantic frames and data sources from the Organisation for Economic Co-operation and Development (OECD). Our findings demonstrate the effectiveness of frame semantics in improving evidence retrieval, indicating a meaningful advancement in automatic fact-checking capabilities. Finally, we conducted a survey of frames evoked in fact-checked claims, identifying high-impact frames to guide future research.
Automatic Fact-Checking with Frame-Semantics
@@ 1 Introduction
The proliferation of misinformation presents a significant challenge to the modern information ecosystem. Mitigating the spread of fake news has become a critical concern for researchers, policymakers, and technology developers alike. In response, automatic fact-checking has emerged as an important research area, with the goal of reducing the burden on human fact-checkers by automating key steps in the fact-checking pipeline.
@@
Many previous studies on automatic fact-checking have utilized unstructured data, such as fact-checks from trustworthy sources Wang (2017), to perform claim matching Shaar et al. (2020) and use large language models to provide verdicts and explanations Cheung and Lam (2023b); Singhal et al. (2024a); Khaliq et al. (2024). Some works have also studied structured data, such as tabular data from Wikipedia Chen et al. (2020); Aly et al. (2021) or scientific documents Wang et al. (2021); Akhtar et al. (2022). These works have been studied on both simple claims extracted from Wikipedia Bouziane et al. (2021) and real-world claims Wang et al. (2021); Akhtar et al. (2022).
Previous approaches Ye et al. (2023); Karagiannis et al. (2020); Jo et al. (2019) have found success with using SQL to query Wikipedia tables, which have an average of 13 rows per table Chen et al. (2020). color=green]The last sentence “due to a lack of …” is a bit confusing . I suppose we are talking about Wikipedia tables only. If so, then the notion of “datasets” is moot, because there is only one dataset that’s applicable—all tables in Wikipedia. By “due to a lack of ”, did you intend to say that most tables there are small? — the point I was making is that there are no existing large-scale datasets for structured automatic fact checking, which is why no past works have addressed it Due to a lack of existing datasets, no previous work has studied automatic fact-checking on real-world claims using large-scale structured data. To overcome the limitations of previous methods, we introduce a novel pilot dataset of real-world claims extracted from PolitiFact  [^1] https://www.politifact.com/ annotated for large-scale structured data.
Our proposed pilot dataset serves as the foundation for two case studies focused on leveraging frame semantics for automatic fact-checking. Frame semantics Ruppenhofer et al. (2016); Baker et al. (1998) is a linguistic framework that explores how language encodes meaning through structured representations called frames. These frames capture the essential elements and relationships in various situations, enhancing the understanding of context and intent behind statements.
In the first case study, we utilize the Vote semantic frame Arslan et al. (2020), which represents how an Agent (e.g., a politician) interacts with a particular Issue (e.g., a proposed bill) through voting. Each frame includes specific components known as frame elements, which provide additional context about the roles and relationships involved, such as the voting member, the bill at stake, and the outcome of the vote. color=cyan]This sentence “Each frame includes …” is a little problematic. The previous sentence already discussed Agent and Issue, which are frame elements. And this sentence is a bit redundant with the previous paragraph. Some re-ordering, reorganization may help. This case study includes 79 claims along with their corresponding congressional members and bills, with an average of 4,230 votes per congress member extracted from official U.S. voting records.color=green]you mentioned the frame side of the framework, but not the database side of it. Particularly, the 4,230 votes per member come from the voting record database. Without discussing the database and its role in this, a reviewer may get confused by the mentioning of the number.
The second case study broadens the scope by exploring a variety of additional semantic frames and a diverse collection of datasets from the Organisation for Economic Co-operation and Development (OECD), which contain over 400 tables with an average of 596,000 rows and 13 columns. color=green]should we say “OECD-related claims” instead? Also, “For OECD claims, we annotated 68 claims” is strange. Couldn’t we just say “We annotated 68 OECD-related claims”? We annotated 68 OECD-related claims along with their corresponding frames and mapped each claim to the OECD statistics used to fact-check them. color=green]The notion of “evoke” hasn’t been mentioned. color=green]Similar to the last paragraph, you haven’t discussed the role of the OECD data in fact-checking. These case studies help to identify the challenges associated with automated fact-checking using large-scale structured data.
Our findings demonstrate the value of frame semantics in automatic fact-checking by enabling a structured and explainable understanding of claims. For instance, on voting claims, we found that using frame elements extracted from claims, instead of the entire claim itself, improved evidence retrieval, leading to a +2.1-point increase in recall@10 for voting claims. Similarly, on OECD claims, we saw a +7.3-point increase in recall@5 when frame elements were used to identify relevant OECD tables. These improvements illustrate how frame-semantics color=cyan]earlier you only said “frame semantics”, not parsing. can enhance retrieval by guiding systems toward relevant structured data.
To understand which frames are most commonly evoked in factual claims, we surveyed claims fact-checked by PolitiFact and found a heavy skew towards a few specific frames, including Taking_sides, Speech, Change_position_on_a_scale. These insights guided the selection of frames for our case studies and enable future works to study high-impact frames.
@@
  • We proposed a novel paradigm for automatic fact-checking using frame-semantics.
@@
  • We developed a pilot dataset for automatic fact-checking of real-world claims using large-scale structured data. [^2] https://github.com/anon-naacl25/
@@
  • We conducted a novel survey of frames evoked in PolitiFact fact-checks, allowing researchers to target high-impact frames for future studies.
@@
  • We conducted two case studies on the efficacy of frame-semantics in automatic fact-checking and released a public demo. [^3] https://anon-naacl25.github.io/
2 Related Works and Background
@@
Recent advances in automatic fact-checking have been largely driven by the integration of large language models (LLMs) and retrieval-augmented generation (RAG) pipelines. Wang et al. (2024) proposed a unified framework for LLM-based systems which utilizes an internal mechanism to determine which of a selection of LLM-base models should be used to fact-check a particular claim. RAGAR Khaliq et al. (2024) enhances fact-checking by using multimodal inputs and iterative reasoning. Similarly, FactLLaMA Cheung and Lam (2023a) combines pre-trained LLaMA models with external evidence retrieval to verify claims. LLM-Augmenter Peng et al. (2023) integrates external knowledge sources and providing feedback to improve model accuracy. Singhal et al. (2024b) mitigates misinformation generated by RAG pipelines via re-ranking retrieved documents according to a credibility score.
@@
For structured data, many approaches have been studied using fine-tuning methods Zhao et al. (2022); Gu et al. (2022); Jiang et al. (2022) and LLM-based methods Ye et al. (2023); Chen et al. (2021); Cheng et al. (2022) for fact verification. Dater Ye et al. (2023) is a detailed fact verification system which decomposes claims into sub-questions using LLMs to simplify the fact-checking process and is the best-performing model on the TabFact Chen et al. (2020) dataset. In this work, we do not focus on fact verification; however, we choose to utilize LLM-based methods for our fact verification step as they do not require retraining the model when new data is used. color=green]I felt this statement undermines the significance of our work. It focuses on the usage of LLM. But the more important essence of our work is the usage of frames.
2.1 End-to-end Fact-Checking Systems
ClaimBuster Hassan et al. (2017) describes the process of automatic fact-checking as having several core components including a claim matching component, where claims are matched with previous fact checks if they exist, and a knowledge base lookup, which can directly answer questions using internal knowledge, e.g., Wolfram|Alpha. Automatic fact-checking systems, like this work, aim to fulfill the role of the knowledge base. color=cyan]Is that internal or external? More recent studies Guo et al. (2022) have decomposed the process into claim detection, evidence retrieval, and verdict prediction and justification. color=green]Looks strange to have a subsection for a short paragraph. Its connection to the whole section and to the paper’s topic is not clearly explained. — paragraph header looks a bit weirder in my opinion. should we just remove this subsection?
@@ 3 Fact-Checking Paradigm
We break down the task of fact-checking into three key steps: claim understanding, evidence retrieval, and claim-evidence alignment. Figure 1 provides an example of how the system processes the statement “Japan’s life expectancy is the 2nd best in Asia” to illustrate this process.
color=green]Need some better re-organization. the first half of this section has several segments, each with a bold-face title. Then we suddenly start to have subsections for the 2nd half, where each subsection is just a paragraph.
@@ 3.1 Claim Understanding
In the first step, the system focuses on understanding the specifics of the claim. Frame-semantic parsing enables the extraction of structured, predefined representations from claims. By identifying frames and their associated frame elements (FEs), the system gains a comprehensive understanding of the claim, which is critical for downstream tasks like evidence retrieval and verification. For example, in Figure 1, the Occupy_Rank frame is evoked, helping the system understand the claim in terms of the Rank of an Item along a Dimension.
Frame semantics not only facilitate claim understanding but also guide the fact-checking process by linking frame elements to appropriate data sources. While some frame elements, such as the Agent and Issue in the Vote frame, map directly to specific database tables, other frames may require a predicted mapping based on the text. Figure 1 illustrates an example of this predicted mapping.
@@ 3.2 Data Collection and Evidence Retrieval
The availability of trustworthy data is essential for fact-checking. Rather than relying on search engines that can return inconsistent results or be vulnerable to fake news injections Huynh and Hardouin (2023); Horne et al. (2019), our system stores ground-truth data in an internal database. Using data from reliable sources like the OECD and U.S. Congress ensures the integrity of the evidence.
Evidence retrieval is guided by the frame elements identified in the claim understanding phase. These frame elements act as filtering conditions for querying the appropriate tables or data sources. For example, in Figure 1, the Dimension frame element is used to query the relevant table for life expectancy statistics. By focusing on specific spans of the claim, this method minimizes confusion and enhances retrieval accuracy. Additionally, due to the fine-grained control enabled by frame-semantics, our proposed system can support unstructured data by adapting methods like retrieval-augmented generation Lewis et al. (2020) using text chunking techniques to query large text corpora instead of a database.
@@ 3.3 Claim-Evidence Alignment
The final step involves aligning the retrieved evidence with the original claim. Frame elements play a critical role in ensuring that the evidence accurately reflects the claim’s context. By only using specific spans of the claim, the system can align the claim and evidence more precisely. This approach also enhances explainability, as the system can point to the exact text used in the query.
Determining the truthfulness of a claim requires synthesizing retrieved evidence and understanding their connection to the claim as well as the intent behind the claim. For example, in Figure 1, the claim states that Japan has the 2nd best life expectancy, yet the data indicates that it is in fact the best. While this claim is not exactly true, the spirit of the claim is true, i.e., that the life expectancy is extremely high in Japan, and thus the model would predict that it is mostly true.color=yellow]This is an intriguing example. But does our approach tackles such a challenge, do we explain how it does that in principle, and do we have empirical results to demonstrate it? — this consideration is included as part of our prompt, but not explicitly discussed in depth as it would be quite difficult to study and probably be its own paper.
@@ 4.1 Voting Records
PolitiFact is a leading fact-checking organization that assesses the accuracy of public statements using verifiable evidence. Voting records  [^4] https://www.politifact.com/voting-record/ are a substantial portion of PolitiFact’s fact-checks. In this case study, we focus on automatically fact-checking voting-related claims using the Vote frame and official U.S. congressional records. Specifically, we target the Agent and Issue Frame Elements (FEs), which refer to the voting entity and the topic of the vote, respectively.
@@
We compiled a dataset of official U.S. congressional voting records from the Congress Github Repository. [^5] https://github.com/unitedstates/congress The dataset includes 342,466 bills from the 93rd Congress to the 117th Congress, with 7,195,798 individual votes across 22,447 roll calls. Each member of Congress cast an average of 4,230 votes from the 101st to the 117th Congress, and the dataset includes a total of 12,677 unique members of Congress since the 1st Congress.
In their fact-checking process, PolitiFact fact-checkers rely on evidence from official records and verified data to assess the truthfulness of claims. In the context of the Vote frame, claims often reference specific congressional bills. To evaluate our system, we constructed an evaluation dataset by extracting all PolitiFact fact-checks made before April 2022 that involve claims related to the Vote frame and reference at least one congressional bill. After manual verification, we collected 79 fact-checks, along with their corresponding bills, to form the evaluation set.
@@
To verify voting records, mapping the Agent FE to the correct member of Congress is essential. We use SQL queries to match congress members whose names are similar to the words in the Agent FE. In cases of name ambiguity, e.g., when two members share the same name, we default to the more recent member.
This stage presents several challenges. Claims often use nicknames, such as “Sleepy Joe” for Joe Biden or “Meatball Ron” for Ron DeSantis. To address this, we extracted two lists of political nicknames from Wikipedia Wikipedia contributors (2024a, b) to map nicknames to their corresponding congress members. Although not exhaustive, these lists should cover many common cases. Similarly, some members go by shortened or preferred names, such as “Joe” instead of Joseph or “Ted” for Rafael Edward Cruz. We handle this by utilizing a list of congress members’ preferred names, supplemented by common alternatives  [^6] https://github.com/carltonnorthern/nicknames when necessary.
@@
Identifying the correct bill based on the extracted Issue FE is challenging because the Issue FE can refer to various types of information, such as abstract topics (e.g., “gun control”), specific actions or bills (e.g., the “Inflation Reduction Act of 2022”), or outcomes of legislation (e.g., “preventing women from getting abortions”). Additionally, bills often do not include the colloquial terms commonly used to describe them. For example, the “STOP School Violence Act of 2018” may be informally described as expanding access to guns in schools, even though the bill itself does not use this phrasing.
Because of these challenges, keyword-based search is insufficient for accurate evidence retrieval. To address this, we employ an asymmetric semantic similarity model, which allows us to identify bills that are semantically similar to the Issue FE, even when the language used in the claim and the bill differs significantly.
@@
Determining whether a claim is refuted or supported by evidence presents several challenges. First, the system cannot rely solely on the vote (Yea or Nay) and the Position Frame Element (FE) (for or against). Claims may be made without a Position FE, and the relationship between the vote on a bill and the claim’s Position FE can differ. For instance, a Yea vote on a bill may not indicate support for the claimed Issue. Consider the claim, “DeSantis voted against allowing abortions,” in conjunction with a Yea vote on a bill that bans abortion. Here, the vote supports the claim despite the discrepancy between the Position FE (against) and the vote on the bill (for/yea). Second, assessing whether a claim is supported or refuted by a bill vote requires an understanding of the bill itself and its implications. color=yellow]This analysis is valid, but do we discuss solution for “understanding of the bill itself and its implications”. Reading this makes me feel it is doing hands-waving.— LLM is prompted to consider potential impacts of voting a certain way on a given bill
Claim Verification.
Finally, our system integrates the alignments between each bill and claim to perform fact verification over all of the retrieved evidence. We instruct a large language model (LLM) to conduct the final verification using the prompt defined in Appendix D.4. color=green]Not clear to me what “each individual claim-bill alignement” means.
Figure 2: System design of OECD case study. The frame-semantic parser (orange) facilitates the claim understanding process, table matching and querying (blue) facilitate the evidence retrieval process, and the claim verification step (green) facilitates the fact verification.
4.2 OECD-Related Statistics
The OECD provides a wealth of trustworthy, observational data on a diverse set of statistics and serves as a strong focal point to explore a wider range of semantic frames. As a visual aid, we include a system overview in Figure 2.

### v2

# Task-Oriented Automatic Fact-Checking with Frame-Semantics

@@ Abstract
We propose a novel paradigm for automatic fact-checking that leverages frame semantics to enhance the structured understanding of claims and guide the process of fact-checking them. To support this, we introduce a pilot dataset of real-world claims extracted from PolitiFact, specifically annotated for large-scale structured data. This dataset underpins two case studies: the first investigates voting-related claims using the Vote semantic frame, while the second explores various semantic frames based on data sources from the Organisation for Economic Co-operation and Development (OECD). Our findings demonstrate the effectiveness of frame semantics in improving evidence retrieval and explainability for fact-checking. Finally, we conducted a survey of frames evoked in fact-checked claims, identifying high-impact frames to guide future work in this direction.
Task-Oriented Automatic Fact-Checking with Frame-Semantics
@@ 1 Introduction
The proliferation of misinformation poses a critical challenge to the modern information ecosystem, threatening informed decision-making and public trust. Addressing this challenge requires scalable and reliable methods to verify the accuracy of claims—particularly novel ones that human fact-checkers may not have previously encountered. Automatic fact-checking has thus emerged as a crucial research area, aiming to reduce the manual burden of verifying claims by automating key steps in the fact-checking pipeline.
@@
Much of the previous work on automatic fact-checking has focused on unstructured data, such as fact-checks from trustworthy sources, to perform claim matching Shaar et al. (2020) and use large language models to produce fact-check verdicts and explanations Cheung and Lam (2023b); Singal et al. (2024); Khaliq et al. (2024). Structured tabular data from Wikipedia Chen et al. (2020); Aly et al. (2021) and scientific documents Wang et al. (2021); Akhtar et al. (2022) have also been utilized to vet synthetic claims extracted from Wikipedia Bouziane et al. (2021) and real-world claims Wang et al. (2021); Akhtar et al. (2022).
However, these studies only consider data which have already been processed and prepared for easy consumption by readers. This reliance on highly-curated data hinders their ability to fact-check novel claims which do not have such clean evidence readily available. To bridge this gap our work proposes a novel pilot dataset of real-world factual claims related to two high-volume structured databases on U.S. congressional voting records and country statistics.color=yellow]Saying we created the dataset to fill the gap gives me the feeling of we invent the nail. Instead, consider saying the verification of many claims rely on data that only exist in much larger tables. These databases significantly differ from past works in both volume and detail of their tables. For example, Wikipedia tables consist of an average of 13 rows Chen et al. (2020) while our database of country statistics has an average of 596,000 rows across over 400 tables. This scale of data is incompatible with most previous methods, which often rely on large language models (LLMs), as the tables themselves are too large to be analyzed, even for those solutions designed for large tables Nahid and Rafiei (2024).
Fact-checking novel claims requires fine-grained methodology to understand the claim and to map important parts of the claim to their relevant data. Furthermore, transparency and explainability are critical in fact-checking. We identify frame semantics Baker et al. (1998) as a viable solution which addresses each of these points. Frame semantics is a linguistic framework that explores how meaning is encoded through structured representations called frames. These frames capture the important elements and relationships in various situations and enable explainable understanding of claims and extraction of their key parts. The existence of specific frames in a claim also enables fine-grained task-specific handling of claims for fact-checking.
Our proposed pilot dataset serves as the foundation for two case studies that demonstrate how frame semantics can be used to guide specific fact-checking processes and enhance the explainability of automatic fact-checking. In the first case study, we focus on the Vote frame Arslan et al. (2020), which models how an Agent (e.g., a legislator) interacts with an Issue (e.g., a bill) through the act of voting and optionally takes a Position (for or against) on the Issue. These Issue and Agent elements are then used to implement voting-specific fact-checking methods. The second case study follows similar principles on a much more diverse set of frames applied to a high-volume, structured database produced from hundreds of datasets collected from the Organisation for Economic Co-operation and Development (OECD).
To evaluate these case studies, we collected 79 and 68 real-world claims for the voting records and OECD statistics, respectively. Each claim is manually annotated for its particular frames and fact-check verdict retrieved from PolitiFact.
Our case studies demonstrate that frame semantics offers a valuable framework for task-specific fact-checking. Frame-semantics can also offer additional performance benefits for particular tasks. For example, on Vote and OECD-related claims, retrieving relevant evidence by querying with extracted task-specific frame elements outperformed using the entire claim by 14% and 11%, respectively. The performance of particular retrieval methodologies also significantly varied between the two case studies, indicating a need to select different approaches for each specific task to achieve optimal performance. Furthermore, instructing large language models (LLMs) to fact-check claims benefits from different prompts depending on the type of claim. These factors further showcase the benefit of a task-specific approach to automatic fact-checking compared to a suboptimal generalized approach.
Finally, to understand which frames are most commonly evoked in factual claims, we surveyed claims fact-checked by PolitiFact and identified a heavy skew toward a few highly frequent frames. Notably, the frames used in our case studies are among the most frequently evoked frames in PolitiFact fact-checks. This alignment highlights the practical relevance of our approach, as it directly addresses the semantic structures most commonly encountered in professional fact-checking. color=yellow]This paragraph is almost like after-thought. need to present it more naturally and perhaps in better places. — doesn’t really feel like there is a better location.
@@
  • We proposed a novel paradigm for task-oriented automatic fact-checking using frame-semantics.
@@
  • We developed a pilot dataset for automatic fact-checking of real-world claims using high-volume structured data. [^1] https://github.com/anon-sub-repo/fact-checker
@@
  • We conducted a novel survey of frames evoked in PolitiFact fact-checks (Table 2), enabling targeting high-impact frames for future studies.
@@
  • We conducted two case studies on the efficacy of frame-semantics in automatic fact-checking and released a public demo and source code. [^2] https://anon-sub-repo.github.io/
2 Related Work
Large-scale datasets.
Previous studies Aly et al. (2021); Gupta et al. (2020); Chen et al. (2020) often use the term “large-scale datasets” to refer to collections with a large quantity of claims and tables. In these cases, the emphasis is on the breadth of the dataset, with a significant number of claims and corresponding tables, typically featuring a modest number of rows per table. For instance, FEVEROUS Aly et al. (2021) focuses on fact-checking against Wikipedia tables, which average around 13 rows per table, enabling techniques that rely on lightweight parsing and exhaustive search methods to handle the smaller, well-curated tabular structures. In contrast, our work addresses high-volume data, where individual tables can contain hundreds of thousands or even millions of rows, as seen in the OECD database used in our case studies, which averages 596,000 rows per table. This shift in scale introduces unique challenges for fact-checking systems, particularly in evidence retrieval.
Automatic fact-checking.
@@ 2 Related Work
Recent advances in automatic fact-checking have been largely driven by the integration of large language models (LLMs) and retrieval-augmented generation (RAG) pipelines. Wang et al. (2025) proposed a unified framework for LLM-based systems which utilizes an internal mechanism to determine which of a selection of LLM-base models should be used to fact-check a particular claim. RAGAR Khaliq et al. (2024) enhances fact-checking by using multimodal inputs and iterative reasoning. FactLLaMA Cheung and Lam (2023a) combines a pre-trained LLaMA model with external evidence retrieval to fact-check claims. LLM-Augmenter Peng et al. (2023) integrates external knowledge sources and provides feedback to improve model accuracy. Singal et al. (2024) mitigates misinformation generated by RAG pipelines via re-ranking retrieved documents according to a credibility score.
@@
For structured data, past work has studied using fine-tuning methods Zhao et al. (2022); Gu et al. (2022); Jiang et al. (2022) and LLM-based methods Ye et al. (2023); Chen et al. (2021); Cheng et al. (2023) for fact verification. Dater Ye et al. (2023) is a fact-verification system which simplifies the fact-checking process by decomposing claims into sub-questions using LLMs. It is the best-performing model on the TabFact dataset Chen et al. (2020). Our work does not focus on fact verification; however, we choose to utilize LLM-based methods for our fact-verification step as they do not require retraining the model when new data is used.
Applications of frame semantics.
Madabushi (2024) compares search methods for finding textual entailment facts in RAG-based systems, including a frame-based method. The authors found that their frame-based approach significantly improved recall@K compared to other search-based methods. Additionally, many previous studies have used frame semantics, or similar semantic parsing methods, in creating task-oriented systems in other domains Neves Ribeiro et al. (2023); Gupta et al. (2018); Chen et al. (2014).
@@ 3 Fact-Checking Paradigm
We break down the task of fact-checking into three key steps: claim understanding, evidence retrieval, and claim-evidence alignment. Figure 1 shows an example of how our system processes the statement “Japan has the 2nd highest life expectancy.” color=yellow]Make it clear in the figure that “Occupy Rank” is the frame. Also mention it in the figure caption. color=yellow]Why is “life-expectancy” schema listed twice, with slight difference (w/ and w/o measure)? color=yellow]In figure 1, spell out FSP as “Frame Semantic Parsing (FSP)” color=yellow]In Figure 1, the yellow boxes are procedures. Other colors are data/concept. But they are all represented as boxes. In addition to color, can we further differentiate them using something else? color=yellow]Figure 1 and this section may appear inconsistent. “claim understanding”, “evidence retrieval”, “claim-evidence aligenment” is no where to be found in the figure. On the other hand, isn’t not clear table mapping, table querying, FSP, fact verfication belong to which of the 3 steps. It is like we have presented two alternative naming systems or workflows.
@@ 3.1 Claim Understanding
In the first step, the system focuses on understanding the specifics of the claim. Frame-semantic parsing Das et al. (2014) enables the extraction of frames and their frame elements from claims.color=green]can make the concept of FSP more clear. Even the next sentence (“identifying frames and … FEs …”) kind of defines it and can be used here instead. color=green]Need references for FSP By extracting this representation, the system gains a comprehensive understanding of the claim, which is critical for downstream tasks like evidence retrieval and verification. For example, in Figure 1, the Occupy_Rank frame is evoked, helping the system understand the claim in terms of the Rank of an Item along a Dimension.color=green]Be consistent. The frame name in the figure doesn’t use underscore.
Frame-semantics not only facilitates claim understanding but also guide the fact-checking process by linking frame elements to appropriate data sources. While some frame elements, such as the Agent and Issue in the Vote frame, map directly to specific database tables (i.e., congress members and bills), other frames may require a predicted mapping based on the text. Figure 1 illustrates an example of this predicted mapping.color=yellow]we can find ways to articulate direct vs. predicted mapping more clearly. Probably won’t have time before deadline.
@@ 3.2 Data Collection and Evidence Retrieval
The availability of trustworthy data is essential for fact-checking. Our system incorporates hand-selected data from existing reliable sources, such as the OECD and U.S. Congress, into an internal database. This ensures the integrity of our evidence and avoids the vulnerabilities of external sources.
Evidence retrieval is guided by the frame elements identified in the claim understanding phase. These frame elements act as filtering conditions for querying the appropriate tables or data sources. For example, in Figure 1, the Dimension frame element is used to query the relevant table for life expectancy statistics. By focusing on specific spans of the claim, this method minimizes confusion and enhances retrieval accuracy. Additionally, due to the fine-grained control enabled by frame-semantics, our proposed system can support unstructured data by querying large text corpora instead of a database. color=yellow]If we don’t have experiments to support this, we don’t want to discuss it here. A future work or discussion section at the end is the right place.
@@ 3.3 Claim-Evidence Alignment
The final step involves aligning the retrieved evidence with the original claim. Frame elements play a critical role in ensuring that the evidence accurately reflects the claim’s context. By only using specific spans of the claim, the system can align the claim and evidence more precisely. This approach also enhances explainability, as the system can point to the text used for evidence retrieval.
Determining the truthfulness of a claim requires synthesizing retrieved evidence and understanding their connection to the claim as well as the intent behind the claim. For example, in Figure 1, the claim states that Japan has the 2nd best life expectancy, yet the data indicates that it is in fact the best. While this claim is not entirely accurate, its underlying premise holds true—Japan has an exceptionally high life expectancy. As a result, the model would predict it as mostly true. color=yellow]This is an intriguing example. But does our approach tackles such a challenge, do we explain how it does that in principle, and do we have empirical results to demonstrate it? — this consideration is included as part of our prompt, but not explicitly discussed in depth as it would be quite difficult to study and probably be its own paper.
@@ 4 Fact-Checking Case Studies
color=yellow]This is really an important ToDo item. Untify the two case studies in the same terminology framework.
color=yellow]Another important ToDo item to strengthen the paper: incorporate more case studies.
@@ 4.1 Voting Records
PolitiFact is a leading fact-checking organization that assesses the accuracy of public statements using verifiable evidence. Voting records are a significant portion of PolitiFact’s fact-checks. In this case study, we focus on automatically fact-checking voting-related claims using the Vote frame and official U.S. congressional records. Specifically, we target the Agent and Issue Frame Elements (FEs), which refer to the voting entity and the topic of the vote, respectively.
@@
We compiled a dataset of official U.S. congressional voting records from the Congress Github Repository. [^3] https://github.com/unitedstates/congress The dataset includes 342,466 bills from the 93rd Congress to the 117th Congress, with 7,195,798 individual votes across 22,447 roll calls. Each member of Congress cast an average of 4,230 votes from the 101st to the 117th Congress, and the dataset includes a total of 12,677 unique members of Congress since the 1st Congress.color=yellow]We are mentioning 93rd, 101st, and 1st. If we only have votes since 101st Congress, why did we include 93rd Congress or even the 1st Congress? At least mentioning statistics for them becomes less relevant. — this is the data we have, it wouldn’t really make sense to say we only have data from the 101st since we technically have members from before
PolitiFact fact-checkers rely on evidence from official records and verified data to assess the truthfulness of claims. 444 https://www.politifact.com/article/2022/mar/31/politifacts-checklist-thorough-fact-checking/ In the context of the Vote frame, claims often reference specific congressional bills. To evaluate our system, we constructed an evaluation dataset by extracting all PolitiFact fact-checks made before April 2022 that involve claims related to the Vote frame and reference at least one congressional bill. After manual verification, we collected 79 fact-checks, along with their corresponding bills, to form the evaluation set.
@@
To verify voting records, mapping the Agent FE to the correct member of Congress is essential. We use SQL queries to match congress members whose names are similar to the words in the Agent FE (details in Appendix C.4). In cases of name ambiguity we default to the more recent member.
This stage presents several challenges. Claims often use nicknames, such as “Sleepy Joe” for Joe Biden or “Meatball Ron” for Ron DeSantis. To address this, we extracted two lists of political nicknames from Wikipedia Wikipedia contributors (2024a, b) to map nicknames to their corresponding congress members. Similarly, some members go by shortened or preferred names, such as “Joe” instead of Joseph or “Ted” for Rafael Edward Cruz. We handle this by utilizing a list 555 https://github.com/carltonnorthern/nicknames of congress members’ preferred names, supplemented by common alternatives when necessary.
@@
Identifying the correct bill based on the extracted Issue FE is challenging because the Issue FE can refer to various types of information, such as abstract topics (e.g., “gun control”), specific actions or bills (e.g., the “Inflation Reduction Act of 2022”), or outcomes of legislation (e.g., “preventing women from getting abortions”). Additionally, bills often do not include the colloquial terms commonly used to describe them. For example, the “STOP School Violence Act of 2018” may be informally described as expanding access to guns in schools, even though the bill itself does not use this phrasing. Because of these challenges, keyword-based search is insufficient for accurate evidence retrieval. To address this, we employ an asymmetric semantic similarity model, as described in Section 5.4, enabling us to identify bills that are semantically similar to the Issue FE, even when the language used in the claim and the bill differs significantly.
@@
Determining whether a claim is refuted or supported by evidence presents several challenges. First, the system cannot rely solely on the vote (Yea or Nay) and the Position Frame Element (for or against).  [^6] Note that Position FE is extracted from claims by our frame semantic parsing model. However, only Agent and Issue FEs are used for evidence retrieval in voting-related claims. Claims may be made without a Position FE, and the relationship between the vote on a bill and the claim’s Position FE can differ. For instance, a Yea vote on a bill may not indicate support for the claimed Issue. Consider the claim, “DeSantis voted against allowing abortions,” in conjunction with a Yea vote on a bill that bans abortion. Here, the vote supports the claim despite the discrepancy between the Position FE (against) and the vote on the bill (for/yea). Second, assessing whether a claim is supported or refuted by a bill vote requires an understanding of the bill itself and its implications. color=yellow]This analysis is valid, but do we discuss solution for “understanding of the bill itself and its implications”. Reading this makes me feel it is doing hands-waving.— LLM is prompted to consider potential impacts of voting a certain way on a given bill
Fact verification.
Finally, our system integrates the alignments between each bill and claim to perform fact verification over all of the retrieved evidence. We instruct a large language model (LLM) to conduct the final verification using the prompt defined in Appendix E.4.
4.2 OECD Statistics